# Jülich-Bonn-Washington Collaboration

### v2

@@ Abstract
Hadronic resonances emerge from strong interactions encoding the dynamics of quarks and gluons. The structure of these resonances can be probed by virtual photons parameterized in transition form factors. In this study, twelve $N^{*}$ and $\Delta$ transition form factors at the pole are extracted from data with the center-of-mass energy from $\pi N$ threshold to $1.8\,{\rm GeV}$, and the photon virtuality $0\leq Q^{2}/{\rm GeV}^{2}\leq 8$. For the first time, these results are determined from a simultaneous analysis of more than one state, i.e., $\sim 10^{5}$ $\pi N$, $\eta N$, and $K\Lambda$ electroproduction data. In addition, about $5\cdot 10^{4}$ data in the hadronic sector as well as photoproduction serve as boundary conditions. For the $\Delta(1232)$ and $N(1440)$ states our results are in qualitative agreement with previous studies, while the transition form factors at the poles of some higher excited states are estimated for the first time. Realistic uncertainties are determined by further exploring the parameter space.
@@
I Introduction
The spectrum and structure of hadronic resonances encode the dynamics of Quantum Chromodynamics (QCD) at intermediate energies where confinement and chiral symmetry breaking lead to the properties of matter as we know it. See Ref. Mai et al. (2023a) for a recent review. Especially the excited states of the nucleon ($N^{*}$) and $\Delta$ resonances, have been studied for decades. On the experimental side, electromagnetic interactions are clean probes of the structure of hadrons, motivating the study of resonances through meson photoproduction reactions Ireland et al. (2020); Thiel et al. (2022). Electroproduction processes Aznauryan et al. (2013); Mokeev and Carman (2022); Carman et al. (2020); Blin et al. (2021); Hiller Blin et al. (2023) provide additional information with an extra degree of freedom, the photon virtuality $Q^{2}$, serving as an energy scale to probe hadron structure, particularly through the electromagnetic transition form factors (TFFs) between excited and ground state baryons Aznauryan and Burkert (2012); Ramalho and Peña (2024) facilitated by experimental progress Ripani et al. (2003); Aznauryan et al. (2009); Mokeev et al. (2012, 2016). The structure encoded in TFFs allows conclusions on the nature of resonances like the $N(1440)$ Burkert and Roberts (2019); they also determine quark transverse charge densities Carlson and Vanderhaeghen (2008); Tiator et al. (2009); Miller (2010).
@@ I Introduction
II Methodology
We calculate the TFFs of twelve selected $N^{*}$ and $\Delta$ states at the poles, based on the latest coupled-channel analyses Mai et al. (2023b) of $\pi N$, $\eta N$, and $K\Lambda$ electroproduction off the proton within the Jülich-Bonn-Washington (JBW) framework Mai et al. (2021, 2022, 2023b). The input at the photon point and for the hadronic final-state interaction is provided by the Jülich-Bonn dynamical coupled-channel model. This unitary framework is based on the meson exchange picture and has been developed over the last decades, see Refs. Rönchen et al. (2013, 2022); Wang et al. (2022) and references therein. A potential, which is derived from a chiral Lagrangian, is iterated in a Lippmann-Schwinger-like equation to obtain the scattering amplitude. The analytic properties of the approach, a prerequisite for determining resonances as poles on the unphysical Riemann sheet, have been discussed in Ref. Döring et al. (2009), see also Ref. Sadasivan et al. (2022). In contrast to the purely hadronic interactions, the interaction of a (real or virtual) photon with the meson-baryon states is parameterized by energy-dependent polynomials Rönchen et al. (2014) and the $Q^{2}$ dependence is introduced in a separable form that fulfills Siegert’s condition Mai et al. (2021) which is a consequence of gauge invariance.
Figure 1: $\Delta$ transition form factors of this work (colored bands). All available literature results from MAID Tiator et al. (2016) and ANL-Osaka solutions are depicted by empty symbols Kamano (2018).
Figure 2: $N^{*}$ transition form factors of this work (colored bands). The bands represent the uncertainties of the extraction, see text for further explanation. ANL-Osaka solutions are depicted by empty symbols Kamano (2018).
@@ II Methodology
To define the TFFs, we may focus on the reactions $\gamma^{*}N\to\pi N$ and $\pi N\to\pi N$ because TFFs at resonance poles are reaction independent. Elastic pion-nucleon scattering is parameterized with the dimensionless partial waves $\tau^{l\pm}$ indicating orbital angular momentum $l$ and total spin $J=l\pm\nicefrac{{1}}{{2}}$ Rönchen et al. (2013). Pion electroproduction is parameterized via multipoles with definite $\gamma^{*}N$ helicity $h=\frac{1}{2},\frac{3}{2}$, isospin $I$, orbital angular momentum $l$, and total angular momentum $J$, denoted as $\mathcal{A}_{h}^{l\pm,I}$, and $\mathcal{S}_{1/2}^{l\pm,I}$, the so-called helicity amplitudes. Note that the latter corresponds to the longitudinal polarization part, which do not contribute to any observables with $Q^{2}=0$.
@@
III Results
Coming to the results, twenty-six states are found in Ref. Rönchen et al. (2018), but in this work only the twelve with pole mass $\text{Re}(z_{p})\leq 1.8\,{\rm GeV}$ and orbital angular momentum $l\leq 3$ are chosen, corresponding to the energy range and the truncation of the angular momentum in Ref. Mai et al. (2023b), namely $N(1535)$, $N(1650)$, $N(1440)$, $N(1710)$, $N(1720)$, $N(1520)$, $N(1675)$, $N(1680)$, $\Delta(1620)$, $\Delta(1232)$, $\Delta(1600)$, $\Delta(1700)$. The TFFs for all these states are shown in Figs. 1 and 2 and are available in numerical form JBW Interactive Scattering Analysis website () (under development).
There are four fits in Ref. Mai et al. (2023b) with similar $\chi^{2}$ which roughly give the scale of uncertainties associated with the extraction; as mentioned before, these uncertainties are larger than those of other studies owing to an extensive exploration of the parameter space.
@@ III Results
As for $N(1440)$, see Figs. 2 and  3, in all four fits we consistently obtain a zero crossing in Re $\text{Re\,}A_{1/2}$, which is also found in Refs. Obukhovsky et al. (2011); Wilson et al. (2012); Segovia et al. (2015); Aznauryan and Burkert (2018); Obukhovsky et al. (2019); Burkert and Roberts (2019); Chen et al. (2019); Julia-Diaz et al. (2007); Kamano (2018). The zero occurs at $Q^{2}=(0.1\pm 0.04)$ GeV2 which is smaller than in Breit-Wigner determinations, and also smaller than in the ANL-Osaka pole determination Kamano (2018) (zero at $Q^{2}=0.2$ GeV2). Such a zero transition indicates that the core of $N(1440)$ can be explained as a radial excitation of the nucleon Ramalho and Peña (2024). However, the structure of the $N(1440)$ is a complicated problem: the contributions from the meson clouds or two-hadron systems seem not to be negligible Meißner and Durso (1984); Zahed et al. (1984); Obukhovsky et al. (2011); Segovia et al. (2015); in some models including the present one [^1] In the Jülich-Bonn model, the $N(1440)$ is always dynamically generated from the $t$- and $u$- channel hadron-exchange potentials, see Ref. Krehl et al. (2000) without the need for a genuine $s$-channel state for $N(1440)$. This is verified by Ref. Wang et al. (2024)., the three-quark core is even less important than the other components Sekihara (2021); Wang et al. (2024).
Figure 3: $N(1440)$ transition form factors at small $Q^{2}$ of this work (colored bands) in comparison with the ANL-Osaka solution depicted by empty symbols Kamano (2018). The error bars at $Q^{2}=0\,{\rm GeV}^{2}$ depict the uncertainties of the photoproduction solution at the pole from Ref. Rönchen et al. (2018).
@@
Figure 4: Transverse charge density (unpolarized $\rho_{0}^{pN^{*}}$, polarized along x-axis $\rho_{T}^{pN^{*}}$) of the $p\to N(1440)$ transition as a function of the transverse position $b$ in the $xy$-plane. The orange band (red line) depicts the uncertainty band of our determination (the result using the MAID 2007 helicity couplings Drechsel et al. (2007)). The inset shows corresponding coordinate decompositions with light/dark shades representing negative/positive values.
IV Conclusion
To summarize, this letter reports the baryon transition form factors for $N^{*}$ and $\Delta$ states defined at the resonance poles. For the first time, this is achieved by analyzing more than one final state at the same time, for altogether $\sim 10^{5}$ data. Additional constraints are provided by a large amount of pion- and (real) photon-induced single-meson data. Special emphasis is paid to a more extensive exploration of the parameter space of the analysis framework, resulting in realistic uncertainties for the TFFs reflecting the data situation and inherent ambiguities.
@@ IV Conclusion
Future data from JLab and other facilities will allow to extend the present analysis to higher photon virtualities mapping the transition region to perturbative QCD. Such data would also allow for the exploration of higher-lying resonances that could be of hybrid nature manifested in an unusual $Q^{2}$ dependence of their transition form factors. In turn, this study also can provide a guidance for future experiments, pinpointing the kinematic regions which are most relevant and least determined to reduce the error bars on the TFF’s.
Acknowledgements. We thank Hiroyuki Kamano, Toru Sato, Craig Roberts, Christian Fischer and Lothar Tiator for helpful discussions. The authors gratefully acknowledge computing time on the supercomputer JURECA Jülich Supercomputing Centre (2021) at Forschungszentrum Jülich under grant no. “baryonspectro”. This work is supported by the NSFC and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the funds provided to the Sino-German Collaborative Research Center TRR110 “Symmetries and the Emergence of Structure in QCD” (NSFC Grant No. 12070131001, DFG Project-ID 196253076-TRR 110). Further support by the CAS through a President’s International Fellowship Initiative (PIFI) (Grant No. 2018DM0034) is acknowledged. This work is also supported by the MKW NRW under the funding code NW21-024-A and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 491111487. The work of RW and MD is supported by the U.S. Department of Energy grant DE-SC0016582 and Office of Science, Office of Nuclear Physics under contract DE-AC05- 06OR23177 (MD). This work contributes to the aims of the U.S. Department of Energy ExoHad Topical Collaboration, contract DE-SC0023598. The work of TM is supported by the PUTI Q1 Grant from University of Indonesia under contract No. NKB-441/UN2.RST/HKP.05.00/2024. The work of UGM and DR is further supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of the CRC 1639 NuMeriQS – project no. 511713970.

### v3

@@ Abstract
Hadronic resonances emerge from strong interactions encoding the dynamics of quarks and gluons. The structure of these resonances can be probed by virtual photons parametrized in transition form factors. In this study, twelve $N^{*}$ and $\Delta$ transition form factors at the pole are extracted from data with the center-of-mass energy from $\pi N$ threshold to $1.8\,{\rm GeV}$, and the photon virtuality $0\leq Q^{2}/{\rm GeV}^{2}\leq 8$. For the first time, these results are determined from a simultaneous analysis of more than one state, i.e., $\sim 10^{5}$ $\pi N$, $\eta N$, and $K\Lambda$ electroproduction data. In addition, about $5\cdot 10^{4}$ data in the hadronic sector as well as photoproduction serve as boundary conditions. For the $\Delta(1232)$ and $N(1440)$ states our results are in qualitative agreement with previous studies, while the transition form factors at the poles of some higher excited states are estimated for the first time. Realistic uncertainties are determined by further exploring the parameter space.
@@
Introduction–The spectrum and structure of hadronic resonances encode the dynamics of Quantum Chromodynamics (QCD) at intermediate energies where confinement and chiral symmetry breaking lead to the properties of matter as we know it. See Ref. Mai et al. (2023a) for a recent review. Especially the excited states of the nucleon ($N^{*}$) and $\Delta$ resonances, have been studied for decades. On the experimental side, electromagnetic interactions are clean probes of the structure of hadrons, motivating the study of resonances through meson photoproduction reactions Ireland et al. (2020); Thiel et al. (2022). Electroproduction processes Aznauryan et al. (2013); Mokeev and Carman (2022); Carman et al. (2020); Blin et al. (2021); Hiller Blin et al. (2023) provide additional information with an extra degree of freedom, the photon virtuality $Q^{2}$, serving as an energy scale to probe hadron structure, particularly through the electromagnetic transition form factors (TFFs) between excited and ground state baryons Aznauryan and Burkert (2012); Ramalho and Peña (2024) facilitated by experimental progress Ripani et al. (2003); Aznauryan et al. (2009); Mokeev et al. (2012, 2016). The structure encoded in TFFs allows conclusions on the nature of resonances like the $N(1440)$ Burkert and Roberts (2019); they also determine quark transverse charge densities Carlson and Vanderhaeghen (2008); Tiator et al. (2009); Miller (2010), and even provide critical information for the exploration of possible hybrid baryons with the gluon as a constituent A New Experiment Run Group Proposal Submitted to Jefferson Lab PAC44 .
@@
Methodology–We calculate the TFFs of twelve selected $N^{*}$ and $\Delta$ states at the poles, based on the latest coupled-channel analyses Mai et al. (2023b) of $\pi N$, $\eta N$, and $K\Lambda$ electroproduction off the proton within the Jülich-Bonn-Washington (JBW) framework Mai et al. (2021, 2022, 2023b). The input at the photon point and for the hadronic final-state interaction is provided by the Jülich-Bonn dynamical coupled-channel model. This unitary framework is based on the meson exchange picture and has been developed over the last decades, see Refs. Rönchen et al. (2013, 2022); Wang et al. (2022) and references therein. A potential, which is derived from a chiral Lagrangian, is iterated in a Lippmann-Schwinger-like equation to obtain the scattering amplitude. The analytic properties of the approach, a prerequisite for determining resonances as poles on the unphysical Riemann sheet, have been discussed in Ref. Döring et al. (2009), see also Ref. Sadasivan et al. (2022). In contrast to the purely hadronic interactions, the interaction of a (real or virtual) photon with the meson-baryon states is parametrized by energy-dependent polynomials Rönchen et al. (2014) and the $Q^{2}$ dependence is introduced in a separable form that fulfills Siegert’s condition Mai et al. (2021) which is a consequence of gauge invariance. For a very brief summary of the formalism, see the Supplemental Material See the attached file for a brief summary of the formalism for the coupled-channel dynamics, amplified figures, and further discussions on the uncertainties  [^1] The Supplemental Material includes Refs. Schweber (1964); Sterman (1993); Zhang et al. (2022) that are relevant to the “Time-Ordered Perturbation Theory” (TOPT), based on which the Lippmann-Schwinger-like equation of this model has been constructed. .
Figure 1: $\Delta$ transition form factors of this work (colored bands). The orange (purple) band with solid (dashed) lines exhibits the real (imaginary) part. The bands represent the uncertainties of the extraction, see text for further explanation. All available literature results from MAID Tiator et al. (2016) and ANL-Osaka solutions are depicted by empty symbols Kamano (2018). Note that there are no literature results (defined at the pole) for the states other than $\Delta(1232)$.
Figure 2: $N^{*}$ transition form factors of this work (colored bands). See the caption of Fig. 1 for the explanations of the symbols. Note that there are no literature results (defined at the pole) for the states other than $N(1440)$.
@@
To define the TFFs, we may focus on the reactions $\gamma^{*}N\to\pi N$ and $\pi N\to\pi N$ because TFFs at resonance poles are reaction independent. Elastic pion-nucleon scattering is parametrized with the dimensionless partial waves $\tau^{l\pm}$ indicating orbital angular momentum $l$ and total spin $J=l\pm\nicefrac{{1}}{{2}}$ Rönchen et al. (2013). Pion electroproduction is parametrized via multipoles with definite $\gamma^{*}N$ helicity $h=\frac{1}{2},\frac{3}{2}$, isospin $I$, orbital angular momentum $l$, and total angular momentum $J$, denoted as $\mathcal{A}_{h}^{l\pm,I}$, and $\mathcal{S}_{1/2}^{l\pm,I}$, the so-called helicity amplitudes. Note that the latter corresponds to the longitudinal polarization part, which do not contribute to any observables with $Q^{2}=0$.
@@
Results–Coming to the results, twenty-six states are found in Ref. Rönchen et al. (2018), but in this work only the twelve with pole mass $\text{Re}(z_{p})\leq 1.8\,{\rm GeV}$ and orbital angular momentum $l\leq 3$ are chosen, corresponding to the energy range and the truncation of the angular momentum in Ref. Mai et al. (2023b), namely $N(1535)$, $N(1650)$, $N(1440)$, $N(1710)$, $N(1720)$, $N(1520)$, $N(1675)$, $N(1680)$, $\Delta(1620)$, $\Delta(1232)$, $\Delta(1600)$, $\Delta(1700)$. The TFFs for all these states are shown in Figs. 1 and 2 and will be available in numerical form JBW Interactive Scattering Analysis website () (under development). The amplified figures for $\Delta(1232)$ and $N(1440)$ can be found in the Supplemental Material See the attached file for a brief summary of the formalism for the coupled-channel dynamics, amplified figures, and further discussions on the uncertainties .
There are four fits in Ref. Mai et al. (2023b) with similar $\chi^{2}$ which roughly give the scale of uncertainties associated with the extraction; as mentioned before, these uncertainties are larger than those of other studies owing to an extensive exploration of the parameter space. There is a more detailed discussion in the Supplemental Material See the attached file for a brief summary of the formalism for the coupled-channel dynamics, amplified figures, and further discussions on the uncertainties .
@@
As for $N(1440)$, see Figs. 2 and  3, in all four fits we consistently obtain a zero crossing in Re $\text{Re\,}A_{1/2}$, which is also found in Refs. Obukhovsky et al. (2011); Wilson et al. (2012); Segovia et al. (2015); Aznauryan and Burkert (2018); Obukhovsky et al. (2019); Burkert and Roberts (2019); Chen et al. (2019); Julia-Diaz et al. (2007); Kamano (2018). The zero occurs at $Q^{2}=(0.1\pm 0.04)$ GeV2 which is smaller than in Breit-Wigner determinations, and also smaller than in the ANL-Osaka pole determination Kamano (2018) (zero at $Q^{2}=0.2$ GeV2). Such a zero transition indicates that the core of $N(1440)$ can be explained as a radial excitation of the nucleon Ramalho and Peña (2024). However, the structure of the $N(1440)$ is a complicated problem: the contributions from the meson clouds or two-hadron systems seem not to be negligible Meißner and Durso (1984); Zahed et al. (1984); Obukhovsky et al. (2011); Segovia et al. (2015); in some models including the present one, the three-quark core is even less important than the other components Sekihara (2021); Wang et al. (2024). Especially, the $N(1440)$ is always dynamically generated from the $t$- and $u$- channel hadron-exchange potentials, without the need for a genuine $s$-channel state Krehl et al. (2000). This is verified by Ref. Wang et al. (2024). Note that Refs. Li (1991); Li et al. (1992) have pointed out the $S_{1/2}$ of a hybrid $N(1440)$ state should be zero. We do not see strong indications of a vanishing $S_{1/2}$: the ${\rm Re}\,S_{1/2}$ in fit 1 is smaller but still non-zero.
Figure 3: $N(1440)$ transition form factors at small $Q^{2}$ of this work (colored bands) in comparison with the ANL-Osaka solution. The error bars at $Q^{2}=0\,{\rm GeV}^{2}$ depict the uncertainties of the photoproduction solution at the pole from Ref. Rönchen et al. (2018). See the caption of Fig. 1 for the explanations of the other symbols.
@@
Figure 4: Transverse charge density (unpolarized $\rho_{0}^{pN^{*}}$, polarized along x-axis $\rho_{T}^{pN^{*}}$) of the $p\to N(1440)$ transition as a function of the transverse position $b$ in the $xy$-plane. The orange band (thick red line) depicts the uncertainty band of our determination (the result using the MAID 2007 helicity couplings Drechsel et al. (2007)). The inset shows corresponding coordinate decompositions with light/dark shades representing negative/positive values.
Conclusion–To summarize, this letter reports the baryon transition form factors for $N^{*}$ and $\Delta$ states defined at the resonance poles. For the first time, this is achieved by analyzing more than one final state at the same time, for altogether $\sim 10^{5}$ data. Additional constraints are provided by a large amount of pion- and (real) photon-induced single-meson data. Special emphasis is paid to a more extensive exploration of the parameter space of the analysis framework, resulting in realistic uncertainties for the TFFs reflecting the data situation and inherent ambiguities.
@@
Future data from JLab and other facilities will allow to extend the present analysis to higher photon virtualities mapping the transition region to perturbative QCD. Such data would also allow for the exploration of higher-lying resonances that could be of hybrid nature manifested in an unusual $Q^{2}$ dependence of their transition form factors A New Experiment Run Group Proposal Submitted to Jefferson Lab PAC44 . In turn, this study also can provide a guidance for future experiments, pinpointing the kinematic regions which are most relevant and least determined to reduce the error bars on the TFF’s.
Acknowledgements–We thank Hiroyuki Kamano, Toru Sato, Craig Roberts, Christian Fischer and Lothar Tiator for helpful discussions. The authors gratefully acknowledge computing time on the supercomputer JURECA Jülich Supercomputing Centre (2021) at Forschungszentrum Jülich under grant no. “baryonspectro”. This work is supported by the NSFC and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the funds provided to the Sino-German Collaborative Research Center TRR110 “Symmetries and the Emergence of Structure in QCD” (NSFC Grant No. 12070131001, DFG Project-ID 196253076-TRR 110). Further support by the CAS through a President’s International Fellowship Initiative (PIFI) (Grant No. 2024PD0022) is acknowledged. This work is also supported by the MKW NRW under the funding code NW21-024-A and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 491111487. The work of RW and MD is supported by the U.S. Department of Energy grant DE-SC0016582 and Office of Science, Office of Nuclear Physics under contract DE-AC05- 06OR23177 (MD). This work contributes to the aims of the U.S. Department of Energy ExoHad Topical Collaboration, contract DE-SC0023598. The work of TM is supported by the PUTI Q1 Grant from University of Indonesia under contract No. NKB-441/UN2.RST/HKP.05.00/2024. The work of UGM and DR is further supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of the CRC 1639 NuMeriQS – project no. 511713970.